OpenStudy (emilyisanemoji):

Solve For W. - (1/4w+8)+8=3 w=52 w=13/4 w=-12 w=-3/4

OpenStudy (jackthegreatest):

-1/4w - 8 + 8 = 3 -1/4w = 3 multiply both sides by 4 -w = 12 w = ?

OpenStudy (emilyisanemoji):

i still dont get it

OpenStudy (jackthegreatest):

what I did was I distributed the negative

OpenStudy (jackthegreatest):

is that the step you're confused on?